OpenStudy (anonymous):

Particles of mud are thrown from the rim of a rolling wheel. If the forward speed of the wheel is v_0, and the radius of the wheel is b, show that the greatest height above ground the mud can go is: \[b+\frac{v_0^2}{2g}+\frac{gb^2}{2v_0^2}\]\] At what point on the rolling wheel does this mud leave? (Note: it is necessary to assume that \[v_0^2\ge bg\] (my workings in next post)

OpenStudy (anonymous):

 So we have two co-ordinate systems: one polar co-ordinates relative to the centre of the wheel, and another 1D Cartesian on the ground. Relative to centre of wheel: \[r=b\cos(\theta)i-b\sin(\theta)j\] \[\theta=\omega t=\frac{v_0}{b}t\] \[r^.=-v_0\sin(\theta)i-v_0\cos(\theta)j\] Relative to ground: \[r^.=(v_0-v_0\sin(\theta))i-v_0\cos(\theta)j\] So where do I go from here?